please help to solve the following
the sum of the first and the third terms of a geometric progression is 40. the sum of the second and the fourth terms is 120. Determine the first term and the common ratio of the progression.

Answer by mananth(16946)

the terms will be
a,ar,ar^2,ar^3
Condition I
a+ar^2=40
a*(1+r^2)=40
(1+r^2) =40/a
condition II
ar+ar^3=120
ar(1+r^2)=120
ar(40/a) =120
40r =120
r=3
plug r to get a